OpenStudy (anonymous):

Find the percent composition of a compound containing tin and chloride if 18.35 g of the compound contains 5.74 g pf tin.

OpenStudy (chmvijay):

18.35 gram of compound contains 5.74 gram of tin 100 gram of compound contains = 5.74*100 /18.35 = 31.28 % and 100- 31.28 = 68.72 % chlorine